# The formation of a soliton gas condensate for the focusing Nonlinear Schrödinger equation

### v1

@@ Appendix A Technical Results
  4. 4.
    $\int_{\eta_{1}}\frac{1}{R_{+}(z)}{\rm d}z\in{\mathbb{R}}_{-}$
  5. 5.
    $\int_{\gamma}\frac{1}{R(z)}{\rm d}z\in i{\mathbb{R}}_{-}$
@@
Properties 1.-3. follows from direct computations. Regarding 4., we rewrite the integral as
$\int_{A}^{-{\overline{A}}}\frac{1}{R(s)}{\rm d}s$ $=-\int_{-a}^{a}\frac{1}{R(x+ib)}{\rm d}x=-\left(\int_{-a}^{0}\frac{1}{R(x+ib)}{\rm d}x+\int_{0}^{a}\frac{1}{R(x+ib)}{\rm d}x\right)$ (A.2)
=$\stackrel{{\scriptstyle\text{2.-3.}}}{{=}}-\left({\overline{\int_{0}^{a}\frac{1}{R(x+ib)}{\rm d}x}}+\int_{0}^{a}\frac{1}{R(x+ib)}{\rm d}x\right)\,.$ (A.3)
We must use some geometry to prove that the equation integral is negative. First, we rewrite it in terms of the angles
$\varphi_{1}=\arg(x+ib-A)=-\pi\,,\quad\varphi_{2}=\arg(x+ib+{\overline{A}})=0\,,\quad\varphi_{3}=\arg(x+ib+A)\,,\quad\varphi_{4}=\arg(x+ib-{\overline{A}})\,,$ (A.4)
as
$\Im\left(\int_{0}^{a}\frac{1}{|x-a|^{\frac{1}{2}}|x-a+2ib|^{\frac{1}{2}}|x+a|^{\frac{1}{2}}|x+a+2ib|^{\frac{1}{2}}e^{i\frac{\varphi_{3}+\varphi_{4}}{2}}}{\rm d}x\right)\,.$ (A.5)
We notice that the angles $\varphi_{3}\in\left(\arctan\left(\frac{b}{a}\right),\arctan\left(\frac{2b}{a}\right)\right)$ and $\varphi_{4}\in\left(\frac{\pi}{2},\pi-\arctan\left(\frac{2b}{a}\right)\right)$; therefore, since $a,b>0$
$\frac{\varphi_{3}+\varphi_{4}}{2}\in\left(\frac{\pi}{2},\pi\right)\,,$ (A.6)
which implies 4. With analogue computations, one can show 5. ∎
@@
$\Re\left(g(z)\right)=-\Re\left(g({\overline{z}})\right),\quad z\in\mathbb{C}^{+}.$ (A.7)
@@
$g(z)=\frac{1}{2}-\frac{2c}{\tau}\int_{A}^{z}\frac{{\rm d}s}{R(s)},\quad z\in\mathbb{C}^{+}$ (A.8)
$g(z)=-\frac{1}{2}-\frac{2c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)},\quad z\in\mathbb{C}^{-}.$ (A.10)
Let $z\in\mathbb{C}^{+}$, then
@@
Let the function $g(z)$ be the unique solution to RHP 3.3. For $z\in\eta_{1}$, we have the representation
$g_{+}(z)-g_{-}(z)=\frac{4c}{\tau}\int_{A}^{z}\frac{1}{R_{-}(s)}{\rm d}s.$ (A.13)
The quantity $g_{+}-g_{-}$ has an analytic extension to $\mathbb{C}_{+}\setminus\left(\eta_{1}\cup\gamma\right)$. We define $u(z)$ to be that analytic extension,
$u(z)=\frac{4c}{\tau}\int_{A}^{z}\frac{1}{R(s)}{\rm d}s,\ \qquad z\in\mathbb{C}_{+}\setminus\left(\eta_{1}\cup\gamma\right),$ (A.14)
where the path of integration is chosen to avoid $\eta_{1}\cup\eta_{2}\cup\gamma$, so that
$u_{-}(z)=g_{+}(z)-g_{-}(z),\ \ z\in\eta_{1}.$ (A.15)
There exists a contour $\eta_{3}$ connecting $A$ with $-{\overline{A}}$ such that
$u(z)\in i\mathbb{R},\qquad z\in\eta_{3}.$ (A.16)
The contour $\eta_{3}$ lies above $\eta_{1}$ and we take it to be oriented from $A$ to $-{\overline{A}}$ (like $\eta_{1}$).
Similarly, for $z\in\eta_{2}$, we have
$g_{+}(z)-g_{-}(z)=\frac{4c}{\tau}\int_{A}^{-\overline{A}}\frac{1}{R_{-}(s)}{\rm d}s+\frac{4c}{\tau}\int_{-A}^{z}\frac{1}{R_{-}(s)}{\rm d}s.$ (A.17)
@@
where again the path is chosen to avoid $\eta_{1}\cup\eta_{2}\cup\gamma$. Note that
$\tilde{u}_{-}(z)=\frac{4c}{\tau}\int_{A}^{-\overline{A}}\frac{1}{R_{-}(s)}{\rm d}s+\frac{4c}{\tau}\int_{-A}^{z}\frac{1}{R_{-}(s)}{\rm d}s,\ \qquad z\in\eta_{2},$ (A.19)
so that
$\tilde{u}_{-}(z)=g_{+}(z)-g_{-}(z),\ \ z\in\eta_{2}.$ (A.20)
There exists a contour $\eta_{4}$ connecting $-A$ with ${\overline{A}}$ such that
$\tilde{u}(z)\in i\mathbb{R},\qquad z\in\eta_{4}.$ (A.21)
The contour $\eta_{4}$ lies below $\eta_{2}$ and we take it to be oriented from $-A$ to ${\overline{A}}$ (like $\eta_{2}$).
Furthermore, the following hold
$\frac{2c}{\tau}\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s=1,\quad\frac{2c}{\tau}\int_{A}^{-{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s=\frac{1}{\tau},\quad-\frac{2c}{\tau}\int_{A}^{\infty}\frac{1}{R(s)}{\rm d}s=\frac{1}{\tau}-1\,.$ (A.22)
@@
First, we verify (A.13), for $z\in\eta_{1}$
@@
where we used the definition of $g(z)$ (3.12). As we did for the function $g(z)$, the extension in the upper-half plane of $u(z)=g_{+}(z)-g_{-}(z)$ for $z\in\eta_{1}$ must have an extra contour of non-analyticity and we impose that such contour is exactly $\gamma$. Local analysis of $u(z)$ in a neighborhood of $z=A$ shows that there exists a single arc emerging from $A$ along which $\Re(u(z))\in i\mathbb{R}$. For a formal discussion, see [34, P.31], or [23, Chapter 3]. Now this arc can be extended maximally, and there are only a finite number of possibilities which we consider in what follows (using either [34, Chapters 3 and 4] or [23, Theorem 3.5, p. 37] ), eventually arriving at the conclusion that it must terminate at $-\overline{A}$. Note that properly speaking we should extend this function to the Riemann surface $\mathcal{R}$, where we would be considering level trajectories of a quadratic differential.
Following [23, Theorem 3.5, p. 37], the maximally extended arc could either diverge to $\infty$, or remain in a compact subset of $\mathbb{C}$. If it remains in a compact subset of the plane it must terminate somewhere, and the only other possibilities are the branch points $-A$, $\overline{A}$, and $-\overline{A}$.
To show that it connects $A$ with $-{\overline{A}}$, we rule out the other possibilities, i.e. we prove that $\eta_{3}$ cannot connect $A$ with $\infty$ or $-A$ or ${\overline{A}}$.
If the contour $\eta_{3}$ would connect $A$ with ${\overline{A}}$, then the integral
$-\frac{2c}{\tau}\int_{A}^{{\overline{A}}}\frac{1}{R(s)}{\rm d}s\,,$
should be purely imaginary. By deformation, one can rewrite the previous integral as
$-\frac{2c}{\tau}\int_{A}^{{\overline{A}}}\frac{1}{R(s)}{\rm d}s=-\frac{2c}{\tau}\left(\int_{A}^{-{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s+\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s+\int_{-A}^{{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s\right)\,.$ (A.24)
Applying Proposition 3.6, we deduce that
$\int_{A}^{-{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s=-\int_{-A}^{{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s\,.$ (A.25)
Therefore, by (3.16),
$-\frac{2c}{\tau}\int_{A}^{{\overline{A}}}\frac{1}{R(s)}{\rm d}s=-\frac{1}{2}\,.$
So, $\eta_{3}$ cannot connect $A$ with ${\overline{A}}$.
If $\eta_{3}$ would connect $A$ with $-A$ then
$-\frac{2c}{\tau}\int_{A}^{-A}\frac{1}{R(s)}{\rm d}s\,,$ (A.26)
should be purely imaginary. By deformation of the contour, we obtain the following equality
$-\frac{2c}{\tau}\int_{A}^{-A}\frac{1}{R(s)}{\rm d}s=-\frac{2c}{\tau}\left(\int_{A}^{-{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s+\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s\right)\,,$ (A.27)
therefore, by applying Proposition 3.6 and (3.16) we deduce that
$-\frac{2c}{\tau}\int_{A}^{-A}\frac{1}{R(s)}{\rm d}s=-\frac{1}{2\tau}-\frac{1}{2}\,,$ (A.28)
which has a non-trivial imaginary part, so $\eta_{3}$ cannot connect $A$ with $-A$.
Finally, if $\eta_{3}$ would connect $A$ with $\infty$ then
$-\frac{2c}{\tau}\int_{A}^{\infty}\frac{1}{R(s)}{\rm d}s\,,$ (A.29)
would be purely imaginary. By Proposition 3.6 and contour deformation, the previous integral is equivalent to
$-\frac{2c}{\tau}\int_{A}^{\infty}\frac{1}{R(s)}{\rm d}s=\frac{c}{\tau}\left(\int_{A}^{\infty}\frac{1}{R(s)}{\rm d}s+\int_{\infty}^{-A}\frac{1}{R(s)}{\rm d}s\right)=\frac{c}{\tau}\left(\int_{\eta_{1}}\frac{1}{R_{-}(s)}{\rm d}s+\int_{\gamma}\frac{1}{R(s)}{\rm d}s\right)=\frac{1}{4\tau}-\frac{1}{4}\,,$ (A.30)
where in the last equality we used (A.28). Therefore, we showed that $\eta_{3}$ must connect $A$ with $-{\overline{A}}$ and the equalities (A.22). We now prove that the contour $\eta_{3}$ is always above $\eta_{1}$. We notice that in this region, we are extending the function
$u_{-}(z)=\frac{4c}{\tau}\int_{A}^{z}\frac{1}{R_{-}(s)}{\rm d}s\,\quad z\in\eta_{1}.$
As a consequence of Proposition 3.6 and equation (3.16)
$\int_{A}^{x_{0}+ib}\frac{1}{R_{-}(s)}{\rm d}s={\overline{\int_{-x_{0}+ib}^{-{\overline{A}}}\frac{1}{R_{-}(s)}{\rm d}s}}\,,\quad\frac{2c}{\tau}\in i{\mathbb{R}}_{+}$ (A.31)
for all $x_{0}\in(0,a)$. The previous equality implies that
$\int_{\eta_{1}}\frac{1}{R_{-}(s)}{\rm d}s\in{\mathbb{R}}.$ (A.32)
Proceeding as in the previous proof, see (A.5), we can show that
$\Re\left(\int_{A}^{x_{0}+ib}\frac{1}{R_{-}(s)}{\rm d}s\right)>0\,,\quad\Im\left(\int_{A}^{x_{0}+ib}\frac{1}{R_{-}(s)}{\rm d}s\right)>0$ (A.33)
for all $x_{0}\in(0,a)$, so $\Im(u(x_{0}+ib))>0,\,\Re(u(x_{0}+ib))<0$
Consider now a point $z_{0}=x_{0}+iy_{0}$ for $x_{0}\in(0,a)$ and $y_{0}>b$, we claim that
$\Im\left(\int_{x_{0}+ib}^{x_{0}+iy_{0}}\frac{1}{R(s)}{\rm d}s\right)<0$ (A.34)
and it’s decreasing as a function of $y_{0}$.
First, for $y_{0}\in(b,+\infty)$, we rewrite the previous integral in terms of the angles
$\varphi_{1}=\arg(x_{0}+iy_{0}-A)\,,\quad\varphi_{2}=\arg(x_{0}+iy_{0}+{\overline{A}})\,,\quad\varphi_{3}=\arg(x_{0}+iy_{0}+A)\,,\quad\varphi_{4}=\arg(x_{0}+iy_{0}-{\overline{A}})\,,$ (A.35)
as
$\Re\left(\int_{b}^{y_{0}}\frac{e^{-i\frac{\varphi_{1}+\varphi_{2}+\varphi_{3}+\varphi_{4}}{2}}}{|x_{0}+iy-a-ib|^{\frac{1}{2}}|x_{0}+iy-a+ib|^{\frac{1}{2}}|x_{0}+iy+a-ib|^{\frac{1}{2}}|x_{0}+iy+a+ib|^{\frac{1}{2}}}{\rm d}y\right)\,.$ (A.36)
We notice that $\Phi=\varphi_{1}+\varphi_{2}+\varphi_{3}+\varphi_{4}$ is a continuous function, which has its minimum value when $y_{0}=b$, where $\Phi=\pi$ and its maximum when $y_{0}=+\infty$, where $\Phi=2\pi$. Therefore, $\cos\left(\frac{\Phi}{2}\right)\leq 0$ and it is equal to zero only when $y_{0}=b$. This implies (A.34) and that the integral is a decreasing function of $y_{0}$.
To conclude, let $x_{0}\in(0,a)$ and consider the integral
$\int_{A}^{x_{0}+iy_{0}}\frac{1}{R(s)}{\rm d}s=\int_{A}^{x_{0}+ib}\frac{1}{R(s)}{\rm d}s+\int_{x_{0}+ib}^{x_{0}+iy_{0}}\frac{1}{R(s)}{\rm d}s\,.$ (A.37)
From the previous proof, we know that the first integral has positive imaginary part, and the second one has negative imaginary part, which is a decreasing function of $y_{0}$. Therefore, for any $x_{0}\in(0,b)$ there exists a value of ${\widetilde{y}}(x_{0})$ such that
$\Im\left(\int_{A}^{x_{0}+ib}\frac{1}{R(s)}{\rm d}s\right)=-\Im\left(\int_{x_{0}+ib}^{x_{0}+i{\widetilde{y}}}\frac{1}{R(s)}{\rm d}s\right)\,,$
and
$\begin{cases}\Im\left(\int_{A}^{x_{0}+iy_{0}}\frac{1}{R(s)}{\rm d}s\right)>0\quad&b<y_{0}<{\widetilde{y}}(x_{0})\\ \Im\left(\int_{A}^{x_{0}+iy_{0}}\frac{1}{R(s)}{\rm d}s\right)<0\quad&y_{0}>{\widetilde{y}}(x_{0})\end{cases}$ (A.38)
This implies that
$\begin{cases}\Re(u(x_{0}+iy_{0}))<0\quad x_{0}\in(0,a)\,,b<y_{0}<{\widetilde{y}}(x_{0})\\ \Re(u(x_{0}+iy_{0}))>0\quad x_{0}\in(0,a)\,,y_{0}>{\widetilde{y}}(x_{0})\end{cases}\,.$ (A.39)
By applying Proposition 3.6, we deduce an analogous result for $x_{0}\in(-a,0)$; therefore, we show that $\eta_{3}$ can be rewritten as
$\eta_{3}=\left\{z\in{\mathbb{C}}\;\big{|}\;z=x_{0}+i{\widetilde{y}}(x_{0}),\;x_{0}\in(-a,a)\right\}\,,$
and it lies above $\eta_{1}$. ∎
@@
  • Since $g(\infty)\in i{\mathbb{R}}$, then $\Re\left(g(z)\right)\to 0$ as $z\to\infty$.
@@
Let $A(z),{\widetilde{A}}(z)$ be the solutions of RHP 2.3 and RHP 2.4, respectively, then
$A(z)=\left(I_{2}+\mathcal{O}\left(\frac{1}{N\left(1+|z|\right)}\right)\right){\widetilde{A}}(z)\,,$ (A.50)
@@
Consider the function ${\mathcal{E}}_{A}(z)=A(z)\tilde{A}^{-1}(z)$ which is analytic on $\mathbb{C}\setminus\Gamma_{1}\cup\Gamma_{2}$. Its jump on $\Gamma_{1}$ equals
$J_{{\mathcal{E}}_{A}}(z)=\tilde{A}_{-}(z)J_{A}(z)J_{\tilde{A}}^{-1}(z)\tilde{A}^{-1}(z)=\tilde{A}_{-}\left(\begin{array}[]{cc}1&0\\ \sum_{j=1}^{N}\frac{c_{j}e^{2i\theta(\lambda_{j})}}{z-\lambda_{j}}+N\int_{\eta_{1}}\frac{h(\lambda)\rho(\lambda)e^{2i\theta(\lambda)}{\rm d}\lambda}{z-\lambda}&1\end{array}\right)\left(\tilde{A}_{-}\right)^{-1}$ (A.53)
$=I_{2}+\tilde{A}_{-}\begin{pmatrix}0&0\\ \frac{q_{1}(z;N)}{N}&0\end{pmatrix}\left(\tilde{A}_{-}\right)^{-1}$ (A.54)
@@
The jump for ${\mathcal{E}}_{A}$ on $\Gamma_{2}$ is similarly evaluated, and one finds that
$J_{{\mathcal{E}}_{A}}(z)=I_{2}+\tilde{A}_{-}(z)\begin{pmatrix}0&\frac{q_{2}(z)}{N}\\ 0&0\end{pmatrix}\left(\tilde{A}_{-}\right)^{-1}$ (A.56)
@@
Since both $A$ and $\tilde{A}$ solve Riemann-Hilbert problems on the same collection of contours, the quantity ${\mathcal{E}}_{A}$ also satisfies a Riemann-Hilbert problem:
@@
Find a $2\times 2$ matrix valued function ${\mathcal{E}}_{A}(z)$ that is analytic in $\mathbb{C}\setminus\left(\Gamma_{1}\cup\Gamma_{2}\right)$, satisfies the normalization condition ${\mathcal{E}}_{A}(z)=I+\mathcal{O}\left(\frac{1}{z}\right)$ as $z\to\infty$, has smooth boundary values on the $+$ and -- side of the contours $\Gamma_{1}$ and $\Gamma_{2}$, and satisfies the jump relation
$\left({\mathcal{E}}_{A}\right)_{+}(z)=\left({\mathcal{E}}_{A}\right)_{-}(z)J_{{\mathcal{E}}_{A}}(z)\ ,\ z\in\Gamma_{1}\cup\Gamma_{2}\ .$ (A.59)
We will prove that this Riemann-Hilbert problem is a so-called "small norm Riemann-Hilbert problem", in the sense that the jump matrix $J_{{\mathcal{E}}_{A}}$ is uniformly close to the trivial jump matrix $I_{2}$. We focus on the quantity $q_{1}(z;N)$, which we will show is analytic in an annular region surrounding the contour $\Gamma_{1}$ (but bounded away from the interval $\eta_{1}$), where it satisfies
@@
Next, we return to $J_{{\mathcal{E}}_{A}}(z)$ on $\Gamma_{1}$ (see formula (A.54)). At this point, it is convenient to assume that the contour $\Gamma_{1}$ encircles not only the interval $\eta_{1}$, but also the collection of contours $\eta_{3},\omega_{3}$, and $\omega_{4}$ as well (see Figure 4.4). This assumption simplifies the representation of $\tilde{A}_{-}$ which we use to estimate the jump matrix $J_{{\mathcal{E}}_{A}}(z)$. Indeed, then for $z\in\Gamma_{1}$ we have the following representation for $\tilde{A}$:
$\tilde{A}(z)=e^{-\left(\ln(N)g(\infty)+y(z)\right)\sigma_{3}}{\mathcal{E}}(z)\Lambda(z)e^{\left(\ln(N)g(z)+y(z)\right)\sigma_{3}}\begin{pmatrix}1&0\\ -N\int_{\eta_{1}}\frac{h(\lambda)\rho(\lambda)}{\lambda-z}e^{2i\theta(\lambda)}{\rm d}\lambda&1\end{pmatrix}$ (A.67)
@@
$\,e^{-\left(\ln(N)g(\infty)+y(z)\right)\sigma_{3}}{\mathcal{E}}(z)\Lambda(z)e^{y(z)\sigma_{3}}\left(\begin{array}[]{cc}0&0\\ \frac{q_{1}(z)e^{-2\ln(N)g(z)\sigma_{3}}}{N}&0\end{array}\right)e^{-y(z)\sigma_{3}}\left({\mathcal{E}}(z)\Lambda(z)\right)^{-1}e^{\left(\ln(N)g(\infty)+y(z)\right)\sigma_{3}}.$ (A.70)
@@
    $g(\infty)\in i{\mathbb{R}}$ and therefore $e^{\pm\ln(N)g(\infty)}$ is bounded in $N$.
@@
$J_{{\mathcal{E}}_{A}}(z)=I_{2}+\mathcal{O}\left(\frac{1}{N}\right),$ (A.72)
where the error term is uniform for all $z$ in an annular region surrounding $\Gamma_{1}$. The analysis of $J_{{\mathcal{E}}_{A}}(z)$ on $\Gamma_{2}\subset\mathbb{C}^{-}$ is entirely similar, and we leave the reader to verify that
$J_{{\mathcal{E}}_{A}}(z)=I_{2}+\mathcal{O}\left(\frac{1}{N}\right),$ (A.73)

### v2

@@ Appendix A Technical Results
Furthermore,
$\int_{\eta_{1}}\frac{1}{R_{+}(z)}{\rm d}z=-\frac{K(\cos(\theta))}{|A|}\,,\quad\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s=-i\frac{K(\sin(\theta))}{|A|}\,,$ (A.2)
where $\theta=\arg(A)$, and $K(k)$ is the complete elliptic integral of the first kind [13, Chap. 19].
@@
Properties 1.-3. follows from direct computations. Regarding (A.2), we define $m=\tan^{2}\left(\frac{\theta}{2}\right)$ and we perform the following change of variables
$z=\frac{i\sqrt{m}w+1}{i\sqrt{m}w-1}|A|\,,$ (A.3)
obtaining that
$\int_{A}^{-{\overline{A}}}\frac{{\rm d}z}{R(z)}=-\frac{i}{2|A|\cos^{2}\left(\frac{\theta}{2}\right)}\int_{-m}^{-1}\frac{{\rm d}w}{i\sqrt{1-w^{2}}\sqrt{1-m^{2}w^{2}}}=-\frac{K(\sqrt{1-m^{2}})}{2|A|\cos^{2}\left(\frac{\theta}{2}\right)}\,,$ (A.4)
$\int_{-{\overline{A}}}^{-A}\frac{{\rm d}z}{R(z)}=-\frac{i}{2|A|\cos^{2}\left(\frac{\theta}{2}\right)}\int_{-1}^{1}\frac{{\rm d}w}{i\sqrt{1-w^{2}}\sqrt{1-m^{2}w^{2}}}=-i\frac{K(m)}{|A|\cos^{2}\left(\frac{\theta}{2}\right)}\,.$ (A.5)
Noticing that $\frac{1-m}{1+m}=\cos(\theta)$ , we can apply the Landen’s transform [13, Chap 19.8] to rewrite the previous elliptic integral in terms of $\theta$ as
$K(\sqrt{1-m^{2}})=2\cos^{2}\left(\frac{\theta}{2}\right)K(\cos(\theta))\,,\qquad K(\sin(\theta))=\cos^{-2}\left(\frac{\theta}{2}\right)K(m)\,,$ (A.6)
this concludes the proof.
@@
$\Re\left(g(z)\right)=-\Re\left(g({\overline{z}})\right),\quad z\in\mathbb{C}_{+}.$ (A.7)
@@
$g(z)=\frac{1}{2}-\frac{2c}{\tau}\int_{A}^{z}\frac{{\rm d}s}{R(s)},\quad z\in\mathbb{C}_{+}$ (A.8)
$g(z)=-\frac{1}{2}-\frac{2c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)},\quad z\in\mathbb{C}_{-}.$ (A.10)
Let $z\in\mathbb{C}_{+}$, then
@@
Let the function $g(z)$ be the unique solution to RHP 3.3. Define the function $u(z)$ for $z\in{\mathbb{C}}_{+}\setminus(\eta_{1}\cup\gamma)$ as
$u(z)=-2g(z)+1=\frac{4c}{\tau}\int_{A}^{z}\frac{1}{R(s)}{\rm d}s,\ \qquad z\in\mathbb{C}_{+}\setminus\left(\eta_{1}\cup\gamma\right),$ (A.13)
where the path of integration is chosen to avoid $\eta_{1}\cup\eta_{2}\cup\gamma$. Let $\eta_{3}$ be the arc of the circle of radius $|A|$ centered at 00 connecting $A$ and $-\overline{A}$. We have
$u_{-}(z)=g_{+}(z)-g_{-}(z),\quad z\in\eta_{1}\,,\qquad u(z)\in i\mathbb{R},\quad z\in\eta_{3}$ (A.14)
@@
where again the path is chosen to avoid $\eta_{1}\cup\eta_{2}\cup\gamma$. Let $\eta_{4}$ be the arc of the circle of radius $|A|$ centered at 00 connecting $-A$ and $\overline{A}$. We have
${\widetilde{u}}_{-}(z)=g_{+}(z)-g_{-}(z),\quad z\in\eta_{2}\,,\qquad u(z)\in i\mathbb{R},\quad z\in\eta_{4}$ (A.16)
Clearly $\eta_{4}$ lies below $\eta_{2}$ and we take it to be oriented from $-A$ to ${\overline{A}}$ (like $\eta_{2}$).
@@
First, we verify the first part of (A.14) for $z\in\eta_{1}$
@@
where we used the definition of $g(z)$ (3.21). Setting $\theta=\arg(A)$, we consider the function $h(\phi)$
$h(\phi)=\int_{|A|e^{i\theta}}^{|A|e^{i\phi}}\frac{{\rm d}s}{R(s)}\,,$ (A.18)
which is the restriction, up to an imaginary constant, of $u(z)$ on the circle of radius $|A|$. Consider now $h^{\prime}(\phi)$
$h^{\prime}(\phi)=\frac{i|A|e^{i\phi}}{R(|A|e^{i\phi})}=\frac{\pm i}{\sqrt{2}|A|(\cos(2\phi)-\cos(2\theta))^{1/2}}\,,$ (A.19)
therefore
$h^{\prime}(\phi)\in\begin{cases}i{\mathbb{R}}&\phi\in(-\theta,\theta)\cup(\pi-\theta,\pi+\theta)\\ {\mathbb{R}}&\phi\in(\theta,\pi-\theta)\end{cases}\,.$ (A.20)
By integration, one concludes the proof.
@@
  • Since $g_{\infty}\in i{\mathbb{R}}$, then $\Re\left(g(z)\right)\to 0$ as $z\to\infty$.
@@
Let $A_{N}(z),{\widetilde{A}}(z)$ be the solutions of RHP 2.3 and RHP 2.4, respectively, then
$A_{N}(z)=\left(I+\mathcal{O}\left(\frac{1}{N\left(1+|z|\right)}\right)\right){\widetilde{A}}(z)\,,$ (A.31)
@@
Consider the function ${\mathcal{E}}_{N}(z)=A_{N}(z)\tilde{A}^{-1}(z)$ which is analytic on $\mathbb{C}\setminus\Gamma_{1}\cup\Gamma_{2}$. Its jump on $\Gamma_{1}$ equals
$J_{{\mathcal{E}}_{N}}(z)=\tilde{A}_{-}(z)J_{A_{N}}(z)J_{\tilde{A}}^{-1}(z)\tilde{A}^{-1}(z)=\tilde{A}_{-}\left(\begin{array}[]{cc}1&0\\ \sum_{j=1}^{N}\frac{c_{j}e^{2i\theta(\lambda_{j})}}{z-\lambda_{j}}+N\int_{\eta_{1}}\frac{h(\lambda)\rho(\lambda)e^{2i\theta(\lambda)}{\rm d}\lambda}{z-\lambda}&1\end{array}\right)\left(\tilde{A}_{-}\right)^{-1}$ (A.34)
$=I+\tilde{A}_{-}\begin{pmatrix}0&0\\ \frac{q_{1}(z;N)}{N}&0\end{pmatrix}\left(\tilde{A}_{-}\right)^{-1}$ (A.35)
@@
The jump for ${\mathcal{E}}_{N}$ on $\Gamma_{2}$ is similarly evaluated, and one finds that
$J_{{\mathcal{E}}_{N}}(z)=I+\tilde{A}_{-}(z)\begin{pmatrix}0&\frac{q_{2}(z)}{N}\\ 0&0\end{pmatrix}\left(\tilde{A}_{-}\right)^{-1}$ (A.37)
@@
Since both $A_{N}$ and $\tilde{A}$ solve Riemann-Hilbert problems on the same collection of contours, the quantity ${\mathcal{E}}_{N}$ also satisfies a Riemann-Hilbert problem:
@@
Find a $2\times 2$ matrix valued function ${\mathcal{E}}_{N}(z)$ that is analytic in $\mathbb{C}\setminus\left(\Gamma_{1}\cup\Gamma_{2}\right)$, satisfies the normalization condition ${\mathcal{E}}_{N}(z)=I+\mathcal{O}\left(\frac{1}{z}\right)$ as $z\to\infty$, has smooth boundary values on the $+$ and -- side of the contours $\Gamma_{1}$ and $\Gamma_{2}$, and satisfies the jump relation
$\left({\mathcal{E}}_{N}\right)_{+}(z)=\left({\mathcal{E}}_{N}\right)_{-}(z)J_{{\mathcal{E}}_{N}}(z)\ ,\ z\in\Gamma_{1}\cup\Gamma_{2}\ .$ (A.40)
We will prove that this Riemann-Hilbert problem is a so-called "small norm Riemann-Hilbert problem", in the sense that the jump matrix $J_{{\mathcal{E}}_{N}}$ is uniformly close to the trivial jump matrix $I_{2}$. We focus on the quantity $q_{1}(z;N)$, which we will show is analytic in an annular region surrounding the contour $\Gamma_{1}$ (but bounded away from the interval $\eta_{1}$), where it satisfies
@@
Next, we return to $J_{{\mathcal{E}}_{N}}(z)$ on $\Gamma_{1}$ (see formula (A.35)). At this point, it is convenient to assume that the contour $\Gamma_{1}$ encircles not only the interval $\eta_{1}$, but also the collection of contours $\eta_{3},\omega_{3}$, and $\omega_{4}$ as well (see Figure 4.4). This assumption simplifies the representation of $\tilde{A}_{-}$ which we use to estimate the jump matrix $J_{{\mathcal{E}}_{N}}(z)$. Indeed, then for $z\in\Gamma_{1}$ we have the following representation for $\tilde{A}$:
$\tilde{A}(z)=e^{-\left(\ln(N)g_{\infty}+y(z)\right)\sigma_{3}}{\mathcal{E}}(z)\Lambda(z)e^{\left(\ln(N)g(z)+y(z)\right)\sigma_{3}}\begin{pmatrix}1&0\\ -N\int_{\eta_{1}}\frac{h(\lambda)\rho(\lambda)}{\lambda-z}e^{2i\theta(\lambda)}{\rm d}\lambda&1\end{pmatrix}$ (A.48)
@@
$\,e^{-\left(\ln(N)g_{\infty}+y(z)\right)\sigma_{3}}{\mathcal{E}}(z)\Lambda(z)e^{y(z)\sigma_{3}}\left(\begin{array}[]{cc}0&0\\ \frac{q_{1}(z)e^{-2\ln(N)g(z)\sigma_{3}}}{N}&0\end{array}\right)e^{-y(z)\sigma_{3}}\left({\mathcal{E}}(z)\Lambda(z)\right)^{-1}e^{\left(\ln(N)g_{\infty}+y(z)\right)\sigma_{3}}.$ (A.51)
@@
    $g_{\infty}\in i{\mathbb{R}}$ and therefore $e^{\pm\ln(N)g_{\infty}}$ is bounded in $N$.
@@
$J_{{\mathcal{E}}_{N}}(z)=I_{2}+\mathcal{O}\left(\frac{1}{N}\right),$ (A.53)
where the error term is uniform for all $z$ in an annular region surrounding $\Gamma_{1}$. The analysis of $J_{{\mathcal{E}}_{N}}(z)$ on $\Gamma_{2}\subset\mathbb{C}_{-}$ is entirely similar, and we leave the reader to verify that
$J_{{\mathcal{E}}_{N}}(z)=I_{2}+\mathcal{O}\left(\frac{1}{N}\right),$ (A.54)